OpenStudy (anonymous):

need help!! the instrucions say find the slope of (-5/2, 1/4) and (1/2,1/5) and the answer is 1/2 but i dont know how to do it!! i need to show the work

OpenStudy (barrycarter):

Slope = rise/run or (change in y) divided by (change in x). Since x changes by 3 and y changes by -1/20, the slope is -1/60, so I suspect you made a typo?

OpenStudy (anonymous):

\[Gradient=m=\frac{y_2-y_1}{x_2-x_1}\]\[=\frac{1/4-1/5}{-5/2-1/2}\]\[=\frac{1/20}{-6/2}\]\[=\frac{1/20}{-3}\]\[=\frac{-1}{60}\]